-4x-2y=-12
4x+8y=-24
1
Expert's answer
2011-11-08T08:48:26-0500
At first let's add equations :& -4x-2y+4x+8y=-12-24
Hence 6y=-36 => y=-6
Then using the second equation of the system we obtain& 4x+8(-6)=-24
4x-48=-24
4x=24
Hence x=6
So the answer is x=6, y=-6